OpenStudy (anonymous):

Find the mean proportional (geometric mean) of 15 and 20.

OpenStudy (anonymous):

Ok for this one I don't know what to do

OpenStudy (anonymous):

ok. i have other questions let me try to see if i can get help with this one.

OpenStudy (anonymous):

the geometric mean of two of two numbers is the square root of their product for instance the geometric mean of 2 and 8 is \[\sqrt{2*8}= \sqrt{16}\]

OpenStudy (anonymous):

Can you solve your problem now?

OpenStudy (anonymous):

so \[SO \it would be \sqrt{15*20}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please help. is it?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

as the formula for geometric mean is \[\sqrt[n]{a _{1}a _{2}a _{3}...a _{n}}\]